40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jan 30 21:29:36 2018, awade, Craig, DailyProgress, PMC, Pre mode cleaner installed in North path 
    Reply  Thu Feb 1 15:25:58 2018, awade, Craig, DailyProgress, PMC, Update working schematic of PMC servo card (older card) D980352-A_AWade20180201Mods.pdf
       Reply  Sat Feb 3 18:08:26 2018, awade, DailyProgress, PMC, A few more modifications to the PMC north board D980352-A_AWade20180202Mods.pdf
          Reply  Fri Jul 6 23:43:54 2018, anchal, DailyProgress, PMC, New modifications to the PMC north board LowPrefFilter.filPMCLowPrefFilterOpt.ipynbU2Buffer.filU2Stability.ipynbD980352-A_AWade20180201Mods_--AGupta20180707Mods.pdf
             Reply  Mon Jul 9 16:15:06 2018, rana, DailyProgress, PMC, New modifications to the PMC north board 
                Reply  Thu Jul 12 15:18:26 2018, anchal, DailyProgress, PMC, New modifications to the PMC north board D980352-AGupta20180712Mods.pdf
                   Reply  Sat Jul 21 14:53:16 2018, awade, DailyProgress, PMC, New modifications to the PMC north board 
                      Reply  Wed Aug 29 12:01:54 2018, anchal, Summary, PMC, Summarization of efforts to fix PMC Servo problems D980352-AGupta20180712Mods.pdfPMC_Servo_LISO_Analysis_Vs_Measured.pdf
                         Reply  Thu Aug 30 12:01:37 2018, rana, Summary, PMC, Summarization of efforts to fix PMC Servo problems 
                         Reply  Mon Sep 24 12:53:26 2018, awade and anchal, Summary, PMC, Summarization of efforts to fix PMC Servo problems: external mixer and LP filter 
Message ID: 2211     Entry time: Thu Jul 12 15:18:26 2018     In reply to: 2210     Reply to this: 2212
Author: anchal 
Type: DailyProgress 
Category: PMC 
Subject: New modifications to the PMC north board 

We earlier found that the intended LPF isn't working so we thought of this external LPF idea. So I checked today the PZT input with LCR meter to see if it is in good condition or not. It gave C = 406.4nF, L=60.6 mH and R = 31kOhm 17MOhm. From the spec sheet, the C value looks 20% below the rated value but the spec value has uncertainity of +- 15%, so maybe our PZT is still good.

With the measured values, I calculated again (fitted using LISO) what good value of output resistor would make it closest to 10Hz pole. The value came out to be 37.1k Ohm. I have replaced this output resistor with 39k Ohm now. I'm attaching updated schematic for future reference.

keywords for search: PMC North Driver Board Schematic

Quote:

should probably put back the 30k resistor. It makes a low pass filter with the PZT capacitance (not the cable capacitance). This is to prevent shorting of the HV drive.

 

Attachment 1: D980352-AGupta20180712Mods.pdf  163 kB  Uploaded Thu Jul 12 18:55:55 2018  | Hide | Hide all
D980352-AGupta20180712Mods.pdf
ELOG V3.1.3-